# Safe Interaction via Monte Carlo Linear-Quadratic Games

### v1

@@ Abstract
Safety is critical during human-robot interaction. But — because people are inherently unpredictable — it is often difficult for robots to plan safe behaviors. Instead of relying on our ability to anticipate humans, here we identify robot policies that are robust to unexpected human decisions. We achieve this by formulating human-robot interaction as a zero-sum game, where (in the worst case) the human’s actions directly conflict with the robot’s objective. Solving for the Nash Equilibrium of this game provides robot policies that maximize safety and performance across a wide range of human actions. Existing approaches attempt to find these optimal policies by leveraging Hamilton-Jacobi analysis (which is intractable) or linear-quadratic approximations (which are inexact). By contrast, in this work we propose a computationally efficient and theoretically justified method that converges towards the Nash Equilibrium policy. Our approach (which we call MCLQ) leverages linear-quadratic games to obtain an initial guess at safe robot behavior, and then iteratively refines that guess with a Monte Carlo search. Not only does MCLQ provide real-time safety adjustments, but it also enables the designer to tune how conservative the robot is — preventing the system from focusing on unrealistic human behaviors. Our simulations and user study suggest that this approach advances safety in terms of both computation time and expected performance. See videos of our experiments here: https://youtu.be/KJuHeiWVuWY.
@@ I Introduction
To address this problem, today’s robots recognize that they are often uncertain about the human’s actions. Instead of assuming the human will follow an exact model, these robots search for plans that are safe across a distribution of human behaviors. There are two general approaches here. The first is a precise method: the robot reasons over the distribution of possible human trajectories, and then chooses the optimal control policy that maximizes safety across the distribution [1, 2]. The second approach is based on approximations: the robot simplifies the system dynamics and objectives, and then obtains a closed-form policy that maximizes safety under the simplified setting [3, 4]. Unfortunately, both types of approaches have significant limitations. Theoretically exact solutions are often computationally intractable; e.g., our autonomous vehicle could not apply these methods in real-time to adjust its behaviors. On the other hand, approximate solutions are imprecise, and may result in unsafe behaviors as the system becomes increasingly complex and nonlinear.
Figure 1: Human and drone moving in a shared workspace. Under our proposed MCLQ safety filter, the drone reasons about worst case human actions within designer-specified bounds, and then selects robust behaviors. For instance, here the drone moves across the table to better prevent a collision.
In this paper we propose a safety filter that robots can leverage to tune their real-time behavior and enhance safety around unpredictable humans. Our proposed filter achieves precise and tractable performance by combining both types of prior approaches. Specifically, our insight is that:
@@
We show an example of our method in Figure 1. The robot iteratively approximates the interaction as a linear-quadratic (LQ) system, and finds an action sequence that maximizes robot performance under the worst case human response. In parallel, the robot refines this action sequence using sampling methods that nosily converge towards the Nash Equilibrium of a zero-sum game. Robots that execute the resulting actions are robust to unpredictable humans within designer-specified bounds: even if the person suddenly changes direction, the autonomous drone has planned a safe trajectory.
@@
Identifying Robust Behavior. To maintain safety despite unpredictable humans, we frame interaction as a zero-sum game. Here the robot tries to minimize its cost while realizing an adversarial human might take actions to maximize that cost. The game’s Nash Equilibrium defines a robot policy that is robust to unexpected human behavior; we develop a method (MCLQ) to tractably identify this robust policy.
@@ II Related Work
Related research has introduced a variety of control algorithms to ensure safety during human-robot interaction [2]. Within this larger field, our approach is most connected to game-theoretic controllers. These methods frame the human and robot as two agents with their own objectives, and solve for robot policies that are robust (i.e., maximize performance) while accounting for the human’s decisions.
Hamilton-Jacobi Methods. Game-theoretic works treat human-robot interaction as interconnected system: both agents have a task they are trying to accomplish, and the actions of each agent affect one another. We can formalize these interconnected dynamical games with Hamilton-Jacobi equations [5]. Precisely solving the Hamilton-Jacobi equations provides the optimal, game-theoretic behavior for each agent; e.g., a policy for the autonomous car that avoids collisions with the human. Hamilton-Jacobi analysis has therefore become a gold standard for safe interaction [1, 6, 7, 8]. Unfortunately, most Hamilton-Jacobi equations have no analytical solution, and numerical methods are often intractable — requiring both exponential time and computational memory [9, 4]. As such, within this paper we develop a real-time approximation that converges towards the ideal solution provided by the Hamilton-Jacobi framework.
Linear-Quadratic Approximations. We are not the first to try to approximate the game-theoretic solution. For example, [10, 11, 12] propose neural network models, and [4, 13, 14] outline open-loop methods. One particularly promising approximation recognizes that we can analytically solve the Hamilton-Jacobi equations in a special case: if the system dynamics are linear and the costs are quadratic [5]. For these linear-quadratic (LQ) games the optimal robot policy is closely connected to linear-quadratic regulators [15]. Robots can solve LQ games quickly, providing a tractable method to obtain safe behaviors that match the gold standard of Hamilton-Jacobi analysis. Indeed — even if the system is not linear-quadratic — recent works have tried to apply sequential linear-quadratic (iLQ) approximations [4, 16]. Under these iLQ approaches the robot iteratively linearizes the dynamics and quadraticizes the cost before solving the resulting LQ game to find the human and robot control gains. On the one hand, iLQ provides an increasingly prevalent method to tractably identify safe controllers. But on the other hand, this approximation is fundamentally limited by its reliance on linear-quadratic systems — as the real interaction diverges from the simplification, iLQ falls short. Overall, this gap motivates our work towards methods that combine theoretical exactness and practical implementation.
@@ III Problem Statement
Although our approach extends to $k$ agents, we will describe interactions between one human and one robot. The robot has a task to perform, and to complete this task the robot must reason over the human (e.g., avoid colliding with the human worker). Without loss of generality, we assume the robot has an initial control policy for the task, as well as a nominal model for predicting the human’s behavior. Our approach will introduce a safety filter that modifies the robot’s policy to enhance safety even when the human deviates from the nominal model. To achieve this safety, we consider the worst-case interaction (within designer specified bounds); i.e., we identify which actions the human can take that will have the worst impact on the robot’s performance.
@@
$\begin{gathered}J\left(x^{t},u^{t:T},w^{t:T}\right)=\sum\limits_{\tau=t}^{T}j(x^{\tau},u^{\tau},w^{\tau})+D(x^{T})\\ \text{s.t.}\leavevmode\nobreak\ x^{\tau+1}=f\left(x^{\tau},u^{\tau},w^{\tau}\right)\end{gathered}$ (2)
@@
In practice, the real human will inevitably deviate from the robot’s model. A worst-case human selects actions that maximize the cost in Equation (2). Formally, in this worst-case the human and robot are participating in a two-player zero-sum game [5]: the robot is trying to minimize its cost, whereas the antagonistic human is attempting to maximize that same cost, directly opposing the robot’s objective.
Nash Equilibria. Similar to prior works [3, 17, 18], we have formulated human-robot interaction as a zero-sum game. The advantage of this formulation is — if we can find robot behavior that is safe in the worst case — then we can ensure safety across the board, regardless of what the human does. Let $\pi_{\mathcal{R}}(x)$ be the robot’s filtered policy, and let $\pi_{H}(x)$ be the human’s true policy. In a zero-sum game, the optimal policies for both agents form a Nash Equilibrium [19], where neither agent can unilaterally deviate from their policy without negatively impacting performance:
$\pi_{\mathcal{R}}$ $=\operatorname*{arg\,min}\limits_{u^{0:T}\in\mathcal{U}^{T}}\max\limits_{w^{0:T}\in\mathcal{W}^{T}}J(x^{t},u^{0:T},w^{0:T})$ (5)
$\pi_{\mathcal{H}}$ $=\operatorname*{arg\,max}\limits_{w^{0:T}\in\mathcal{W}^{T}}\min\limits_{u^{0:T}\in\mathcal{U}^{T}}J(x^{t},u^{0:T},w^{0:T})$ (6)
In a zero-sum game, the Nash Equilibrium is unique [19]: there is a single set of (mixed) policies that satisfy Equations (5)–(6). By definition, the robot policy of Equation (5) is maximally robust to changes in the human policy. This policy enables us to maintain safe interactions — even when we do not know what actions the real human will take.
Solving for Nash Equilibria. Accordingly, it is desirable to shift the robot’s initial policy $\hat{\pi}_{\mathcal{R}}$ towards the Nash Equilibrium policy. But how do we solve for this equilibrium? The standard approach in finite horizon, discrete time settings is to apply the Bellman Equation:
@@
Applying Equation (7) in continuous time settings results in the Hamilton-Jacobi-Isaacs (HJI) partial differential equation. However, this formula can only be applied tractably in special cases such as linear-quadratic systems [5]; in general, the Nash Equilibrium policy must be approximated.
Objective. Given that the robot has some initial policy $\hat{\pi}_{\mathcal{R}}$, our goal is to develop a real-time safety filter that adjusts the robot’s policy so that it converges towards the Nash Equilibrium solution. Because we cannot tractably identify this ideal policy, we must develop an approximation that the robot can leverage efficiently in real-time.
@@ IV Monte Carlo Linear-Quadratic Games
In this section we propose our real-time safety filter based on zero-sum games. We recognize that there are special cases where we can find analytical solutions to Equation (7): when the dynamics are linear and the cost is quadratic. Accordingly, we first leverage this linear-quadratic case to obtain an initial guess of the Nash Equilibrium policies (Section IV-A). We next refine that guess by stochastically solving the true optimization problem (Section IV-B). Our overall approach — Monte Carlo Linear-Quadratic Games (MCLQ) — combines rapid LQ solutions with a parallel local search. We theoretically demonstrate that MCLQ improves upon existing LQ baselines, and converges towards the ideal Nash Equilibrium policy (Section IV-C). We conclude with the practical advantages of our MCLQ framework, including a designer-specified safety margin that prevents the robot from becoming overly conservative (Section IV-D).
@@ IV-A Obtaining an Initial Action Trajectory
where $\circ^{\prime}$ represents transposition. Under this LQ approximation the Nash Equilibrium policies exhibit linear state-feedback behavior [5]. These optimal feedback policies are captured by the gain matrices:
@@
To find the gain matrices, we employ the discrete-time algebraic Riccati Equation (DARE, a version of the Bellman Equation) as detailed in [15]. For a particular robot gain matrix $K$, the optimal human gain matrix $L$ can be found in closed form by solving the DARE:
@@
and the terminal condition is $\bm{P}_{\bm{K},\bm{L}(\bm{K})}=\bm{Q}$. The bolded matrices shown in Equation (11) are augmented forms of the LQ matrices from Equations (8) and (9):
$\begin{gathered}\bm{A}\equiv\begin{bmatrix}\bm{0}&\bm{0}\\ \text{diag}(A^{t:(T-1)})&\bm{0}\end{bmatrix}\\ \bm{B}\equiv\begin{bmatrix}\bm{0}&\bm{0}\\ \text{diag}(B^{t:(T-1)})&\bm{0}\end{bmatrix}\\ \bm{D}\equiv\begin{bmatrix}\bm{0}&\bm{0}\\ \text{diag}(D^{t:(T-1)})&\bm{0}\end{bmatrix}\\ \bm{Q}\equiv\text{diag}(Q^{t:T})\\ \bm{R}_{u}\equiv\text{diag}(R_{u}^{t:T}),\bm{R}_{w}\equiv\text{diag}(R_{w}^{t:T})\end{gathered}$ (13)
Once the DARE has converged, the unique human gain matrix $\bm{L}(\bm{K})$ that maximizes the cumulative cost is:
@@
By iteratively recomputing the DARE for both Equation (14) and Equation (15), we converge to feedback policies that form the Nash Equilibrium for zero-sum LQ games. Expanding these matrices provides our initial guess of the optimal action trajectory $\xi(x^{0})=(u^{0:T},w^{0:T})$.
@@ IV-B Refining the Action Trajectory
The action trajectory $\xi(x^{0})=(u^{0:T},w^{0:T})$ we have obtained is optimal if the system is linear-quadratic. But often this is not the case; solutions that assume linear dynamics and quadratic costs may fall apart when faced with nonlinear, realistic interactions, or when the necessary Eigenvalue conditions are not met [15]. We therefore propose to use the solution from Section IV-A as a first pass approximation that we will refine with stochastic gradient descent.
As a reminder, our ultimate goal is to reach the robot’s Nash Equilibrium policy. This policy is defined by Equation (5), which can be treated as a nested optimization problem. In the outer loop the robot proposes a sequence of actions $u^{0:T}$ to minimize its cost, and in the inner loop the antagonistic human responds with actions $w^{0:T}$ that maximize cost. We therefore apply a nested gradient descent algorithm to iteratively improve the initial action trajectory $\xi(x^{0})=(u^{0:T},w^{0:T})$ and optimize Equation (5). More specifically, we perform nested Metropolis-Hastings sampling (i.e., stochastic gradient descent). Other Monte Carlo methods can be leveraged within our general framework; we here selected the Metropolis-Hastings sampler due to its ergodic properties and avoidance of local optima [20].
Given the seed $\xi(x^{0})=(u^{0:T},w^{0:T})$ provided by Section IV-A, this sampler has two parts [21]: an inner loop that updates the antagonistic human actions $\xi_{w}=w^{0:T}$, and an outer loop that improves the robot’s actions $\xi_{u}=u^{0:T}$. An outline of our resulting Monte Carlo Linear-Quadratic Games approach is included in Algorithm 1.
@@
Summary. Taken together, the LQ game in Section IV-A provides a first guess at the robot’s action trajectory $\xi$. This action trajectory is then refined by nested Metropolis-Hastings samplers that update $\xi$ to solve for the true Nash Equilibrium policy in Equation (5). The resulting trajectory $\xi^{\prime}$ contains the sequence of actions a robust robot should execute over the next $T$ timesteps; this robust action sequence can be recomputed online (i.e., at each timestep $t$).
Algorithm 1 Monte Carlo Linear-Quadratic Games (MCLQ)
1:$f,J,\beta,M,N$triangleright Dynamics and Cost Functions
2:procedure LQ Approximation($x$, $f$, $J$)
3:     $\bm{A},\bm{B},\bm{D}\leftarrow\text{Linearize}(f)$
4:     $\bm{Q},\bm{R}_{u},\bm{R}_{w}\leftarrow\text{Quadraticize}(J)$
[…]
6:     $\xi_{u}\leftarrow-\bm{K}\bm{x}$
7:     $\xi_{w}\leftarrow-\bm{L}\bm{x}$
8:     return $\xi_{u}$, $\xi_{w}$
9:end $\triangleright$ Provides initial guess of $\xi_{u}$, $\xi_{w}$
10:procedure Monte Carlo Search($x$, $\xi_{u}$, $\xi_{w}$, $f$, $J$)
11:     $\Delta_{u}\leftarrow 0$, $\Delta_{w}\leftarrow 0$
12:     $J_{0}\leftarrow J(\xi_{u},\xi_{w})$
13:     for $m=1\ldots M$ do
14:         for $n=1\ldots N$ do
15:              $\Delta_{n}\leftarrow\text{Perturb}(\xi_{w},\Delta_{w})$
16:              $J_{w}\leftarrow J(\xi_{u}+\Delta_{u},\xi_{w}+\Delta_{n})$
17:              if $\exp(\beta(J_{w}-J_{0}))>\eta,\leavevmode\nobreak\ \eta\sim U[0,1]$ then
18:                  $J_{0}\leftarrow J_{w}$
19:                  $\Delta_{w}\leftarrow\Delta_{n}$
20:              end if
21:         end $\triangleright$ Updates worst-case human actions
22:         $\Delta_{m}\leftarrow\text{Perturb}(\xi_{u},\Delta_{u})$
23:         $J_{u}\leftarrow J(\xi_{u}+\Delta_{m},\xi_{w}+\Delta_{w})$
24:         if $\exp(\beta(J_{0}-J_{u}))>\eta,\leavevmode\nobreak\ \eta\sim U[0,1]$ then
25:              $J_{0}\leftarrow J_{u}$
26:              $\Delta_{u}\leftarrow\Delta_{m}$
27:         end $\triangleright$ Updates robot response to human
28:     end for
29:     return $\xi_{u}^{\prime}=\xi_{u}+\Delta_{u}$, $\xi_{w}^{\prime}=\xi_{w}+\Delta_{w}$
30:end procedure
IV-C Theoretical Analysis
Here we theoretically compare MCLQ to both the ideal solution (i.e., the Nash Equilibrium) and relevant approximations (i.e., LQ games). We consider two cases. First, when the actual system has linear dynamics $f$ and a quadratic cost $J$, and second, when the system is not linear-quadratic.
@@ IV-C Theoretical Analysis
LQ Systems. For a linear-quadratic system the Bellman Equation is equivalent to the DARE presented in Equation (11) [22]. Hence, the trajectory produced by the DARE is the Nash Equilibrium of the zero-sum game. This means that — for our MCLQ method — the initial trajectory is the global optimum of Equation (5), and Monte Carlo sampling is unable to improve upon this initial guess. Each method therefore converges to the same sequence of robot actions.
Non-LQ Systems. If the system is not linear-quadratic, then MCLQ improves upon the LQ approximation. Consider the case where $\beta\rightarrow\infty$; here our sampler only accepts perturbations $\Delta_{u}$ and $\Delta_{w}$ if they move towards the Nash Equilibrium in Equation (5). In the worst case, the initial trajectory provided by the DARE is a local optima, and MCLQ performs on par with LQ games. In any other case, the robot is not seeded in a local optima, and nested gradient descent moves towards the Nash Equilibrium. Hence, MCLQ matches or outperforms standard LQ approximations.
IV-D Balancing Performance and Safety

### v2

# Safe Interactions via Monte Carlo Linear-Quadratic Games

@@ Abstract
Safety is critical during human-robot interaction. But — because people are inherently unpredictable — it is often difficult for robots to plan safe behaviors. Instead of relying on our ability to anticipate humans, here we identify robot policies that are robust to unexpected human decisions. We achieve this by formulating human-robot interaction as a zero-sum game, where (in the worst case) the human’s actions directly conflict with the robot’s objective. Solving for the Nash Equilibrium of this game provides robot policies that combine safety and performance across a wide range of human actions. Existing approaches attempt to find these optimal policies by leveraging Hamilton-Jacobi analysis (which is intractable) or linear-quadratic approximations (which are inexact). By contrast, in this work we propose a computationally efficient and theoretically supported method that approaches open-loop Nash Equilibrium policies. Our approach (which we call MCLQ) leverages linear-quadratic games to obtain an initial guess at safe robot behavior, and then iteratively refines that guess with a Monte Carlo search. Not only does MCLQ provide real-time safety adjustments, but it also enables the designer to tune how conservative the robot is — preventing the system from focusing on unrealistic human behaviors. Our simulations and user study suggest that this approach advances safety in terms of both computation time and expected performance.
@@ I Introduction
To address this problem, today’s robots recognize that they are often uncertain about the human’s actions. Instead of assuming the human will follow an exact model, these robots search for plans that are safe across a distribution of human behaviors. There are two general approaches here. The first is a precise method: the robot reasons over the distribution of possible human trajectories, and then chooses the optimal control policy that maximizes safety across the distribution [bansal2017hamilton, hsu2023safety]. The second approach is based on approximations: the robot simplifies the system dynamics and objectives, and then obtains a closed-form policy that maximizes safety under the simplified setting [tian2022safety, fridovich2020efficient]. Unfortunately, both types of approaches have significant limitations. Theoretically exact solutions are often computationally intractable; e.g., our autonomous vehicle could not apply these methods in real-time to adjust its behaviors. On the other hand, approximate solutions are imprecise, and may result in unsafe behaviors that are locally optimal but globally unsafe as the system becomes increasingly complex and nonlinear.
Figure 1: Human and drone moving in a shared workspace. Under our proposed method (MCLQ), the drone reasons about worst case human actions within designer-specified bounds, and then selects safe behaviors in response to those actions. For instance, here the drone moves across the table to better prevent a potential collision.
In this paper we propose an action-augmentation method that robots can leverage to tune their real-time behavior and enhance safety around unpredictable humans. Our proposed method achieves precise and tractable performance by combining both types of prior approaches. Our insight is that:
@@
We show an example of our method in Figure 1. The robot iteratively approximates the interaction as a linear-quadratic (LQ) system, and finds an action sequence that maximizes robot performance under the worst case human response. In parallel, the robot refines this action sequence using sampling methods that noisily converge towards the open-loop optimal of a zero-sum game while stochastically escaping local optima. Robots that execute the resulting actions are robust to unpredictable humans within designer-specified bounds: even if the person suddenly changes direction, the autonomous drone has planned a safe trajectory.
@@
Identifying Robust Behavior. To maintain safety despite unpredictable humans, we frame interaction as a zero-sum game. Here the robot tries to minimize its cost while realizing an adversarial human might take actions to maximize that cost. The game’s Nash Equilibrium defines a robot policy that is robust to unexpected human behavior; we develop a method (MCLQ) to tractably approximate this robust policy.
@@ II Related Work
Related research has introduced a variety of control algorithms to ensure safety during human-robot interaction [hsu2023safety]. Within this larger field, our approach is most connected to game-theoretic controllers. These methods frame the human and robot as two agents with their own objectives, and solve for robot policies that are robust (i.e., maximize performance) while accounting for the human’s decisions.
Hamilton-Jacobi Methods. Game-theoretic works treat human-robot interaction as interconnected system: both agents have a task they are trying to accomplish, and the actions of each agent affect one another. We can formalize these interconnected dynamical games with Hamilton-Jacobi equations [bacsar1998dynamic]. Precisely solving the Hamilton-Jacobi equations provides the optimal, game-theoretic behavior for each agent; e.g., a policy for the autonomous car that avoids collisions with the human. Hamilton-Jacobi analysis has therefore become a gold standard for safe interaction [bansal2017hamilton, bansal2020hamilton, fridovich2020confidence, bajcsy2019efficient]. Unfortunately, most Hamilton-Jacobi equations have no analytical solution, and numerical methods are often intractable — requiring both exponential time and computational memory [starr1969nonzero, fridovich2020efficient]. As such, within this paper we develop a real-time approximation that converges towards the ideal solution provided by the Hamilton-Jacobi framework.
Linear-Quadratic Approximations. We are not the first to try to approximate the game-theoretic solution. For example, [jiang2016using, nakamura2021adaptive, nakamura2020causality] propose neural network models, and [fridovich2020efficient, jones2020polynomial, chilan2020optimal] outline open-loop methods. One particularly promising approximation recognizes that we can analytically solve the Hamilton-Jacobi equations in a special case: if the system dynamics are linear and the costs are quadratic [bacsar1998dynamic]. For these linear-quadratic (LQ) games the optimal robot policy is closely connected to linear-quadratic regulators [wu2023learning]. Robots can solve LQ games quickly, providing a tractable method to obtain safe behaviors that match the gold standard of Hamilton-Jacobi analysis. Indeed — even if the system is not linear-quadratic — recent works have tried to apply sequential linear-quadratic (iLQ) approximations [fridovich2020efficient, laine2023computation]. Under these iLQ approaches the robot iteratively linearizes the dynamics and quadraticizes the cost before solving the resulting LQ game to find the human and robot control gains. On the one hand, iLQ provides an increasingly prevalent method to tractably identify safe controllers. But on the other hand, this approximation is fundamentally limited by its reliance on linear-quadratic systems — as the real interaction diverges from the simplification, iLQ falls short. Overall, this gap motivates our work towards methods that combine theoretical exactness and practical implementation.
@@ III Problem Statement
Although our approach extends to $k$ agents, we will describe interactions between one human and one robot. The robot has a task to perform, and to complete this task the robot must reason over the human (e.g., avoid colliding with the human worker). Without loss of generality, we assume the robot has an initial control policy for the task, as well as a nominal model for predicting the human’s behavior. Our approach will adjust the robot’s policy to enhance safety even when the human deviates from the nominal model. To achieve this safety, we consider the worst-case interaction (within designer-specified bounds): i.e., we identify which actions the human can take that will have the worst impact on the robot’s performance.
@@
$\begin{gathered}J\left(x^{t},u^{t:T},w^{t:T}\right)=\sum\limits_{\tau=t}^{T}j(x^{\tau},u^{\tau},w^{\tau})+D(x^{T})\\ \text{s.t.}\penalty 10000\ x^{\tau+1}=f\left(x^{\tau},u^{\tau},w^{\tau}\right)\end{gathered}$ (2)
@@
In practice, the real human will inevitably deviate from the robot’s model. A worst-case human selects actions that maximize the cost in Equation (2). Formally, in this worst-case the human and robot are participating in a two-player zero-sum game [bacsar1998dynamic]: the robot is trying to minimize its cost, whereas the antagonistic human is attempting to maximize that same cost, directly opposing the robot’s objective.
In this work, we assume that the cost function $J$ and state transition function $f$ are sufficiently smooth and have continuous first- and second-derivatives with respect to $x$. We further assume that there exists a norm $\|\circ\|$ such that $\|\nabla_{x}f(x,u)\|\leq 1$, i.e. the system state does not grow unbounded for finite $u$.
Nash Equilibria. Similar to prior works [tian2022safety, riskawareslumbers2023game, wang2020game], we have formulated human-robot interaction as a zero-sum game. The advantage of this formulation is — if we can find robot behavior that is safe in the worst case — then we can ensure safety across the board, regardless of what the human does. Let $\pi_{\mathcal{R}}(x)$ be the robot’s filtered policy, and let $\pi_{H}(x)$ be the human’s true policy. In a zero-sum game, the optimal policies for both agents form a Nash Equilibrium [nash1951non], where neither agent can unilaterally deviate from their policy without negatively impacting performance. Below we describe the open-loop Nash Equilbrium:
$\pi_{\mathcal{R}}$ $=\mathop{\arg\min}\limits_{u^{0:T}\in\mathcal{U}^{T}}\max\limits_{w^{0:T}\in\mathcal{W}^{T}}J(x^{t},u^{0:T},w^{0:T})$ (5)
$\pi_{\mathcal{H}}$ $=\mathop{\arg\max}\limits_{w^{0:T}\in\mathcal{W}^{T}}\min\limits_{u^{0:T}\in\mathcal{U}^{T}}J(x^{t},u^{0:T},w^{0:T})$ (6)
In a zero-sum game, the closed-loop Nash Equilibrium is unique under mixed-strategies [nash1951non]. Although Equation (5) is open-loop, it is more robust to changes in the human policy than policies that do not reason over the human’s actions. This policy enables us to improve safety — even when we do not know what actions the real human will take.
In general, the open-loop NE of Equations (5) and (6) are not guaranteed to exist, and may not be equivalent to the closed-loop NE. In this work, we will assume that these open-loop equilibrium exist. In some zero-sum linear-quadratic games — especially as the number of players increases — the open- and closed-loop solutions converge [liang2025open]. If the open- and closed-loop optima for the system are locally similar, then solving for open-loop equlibria may yield policies that are more robust to human actions than solving for the LQ-approximate closed-loop NE; however, local similarity is not a requirement for our proposed method.
Solving for Nash Equilibria. It is desirable to shift the robot’s initial policy $\hat{\pi}_{\mathcal{R}}$ towards the Nash Equilibrium policy. But how do we solve for this equilibrium? The standard approach in finite horizon, discrete time settings is to apply the Bellman Equation:
@@
However, this formula can only be applied tractably in special cases such as linear-quadratic systems [bacsar1998dynamic]; in general, the Nash Equilibrium policy must be approximated.
Objective. Given that the robot has some initial policy $\hat{\pi}_{\mathcal{R}}$, our goal is to develop a real-time approach that adjusts the robot’s policy so that it converges towards behavior that is robust to deviations from our human model. Because we cannot tractably identify this ideal policy (e.g., the Nash Equilibrium policy), we must develop an approximation that the robot can leverage efficiently in real-time.
@@ IV Monte Carlo Linear-Quadratic Games
In this section we propose our real-time method for safe interactions based on zero-sum games. We recognize that there are special cases where we can find analytical solutions to Equation (7): when the dynamics are linear and the cost is quadratic. Accordingly, we first leverage this linear-quadratic case to obtain an initial guess of the Nash Equilibrium policies (Section IV-A). We next refine that guess by stochastically solving the open-loop optimization problem (Section IV-B). Our overall approach — Monte Carlo Linear-Quadratic Games (MCLQ) — combines rapid LQ solutions with a parallel local search. We theoretically demonstrate that MCLQ improves upon existing LQ baselines, and converges towards a robust robot policy.We conclude with the practical advantages of our MCLQ framework, including a designer-specified safety margin that prevents the robot from becoming overly conservative (Section IV-C).
@@ IV-A Obtaining an Initial Action Trajectory
where $\circ^{\prime}$ represents transposition. We perform this LQ approximation about the most recently calculated trajectory. If this approximation is being performed at $t=0$, then we sample a random inital trajectory. Under this LQ approximation the Nash Equilibrium policies exhibit linear state-feedback behavior [bacsar1998dynamic]. These optimal feedback policies are captured by the gain matrices:
@@
To find the gain matrices, we employ the discrete-time algebraic Riccati Equation (DARE, a version of the Bellman Equation) as detailed in [wu2023learning]. For a particular robot gain matrix $K$, the optimal human gain matrix $L$ can be found in closed form by solving the DARE:
@@
and the terminal condition is $\bm{P}_{\bm{K},\bm{L}(\bm{K})}=\bm{Q}$. The bolded matrices shown in Equation (11) are lower triangular matrices with values along the diagonal set to 0. For details, we direct the reader to Equation (4) of [wu2023learning]. Once the DARE has converged, the unique human gain matrix $\bm{L}(\bm{K})$ that maximizes the cumulative cost is:
@@
By iteratively approximating the system as LQ and recomputing the coupled DAREs for both Equation (13) and Equation (14), we converge to feedback policies that form the Nash Equilibrium for zero-sum LQ games. Expanding these matrices provides our initial guess of the optimal action trajectory $\xi(x^{0})=(u^{0:T},w^{0:T})$. The traditional formulation of linear-quadratic approximations for zero-sum games can violate the Eigenvalue conditions of Equation (11), even when a closed-form NE exists for the underlying non-LQ system. In practice, this can be alleviated by adjusting $R_{u}$ and $R_{w}$ such that $\|R_{u}\|\ll\|R_{w}\|$ [bacsar1998dynamic]. The gain matrices can be readily calculated using open-source software, such as Scipy’s solve_discrete_are [virtanen2020scipy].
@@ IV-B Refining the Action Trajectory
The action trajectory $\xi(x^{0})=(u^{0:T},w^{0:T})$ we have obtained is optimal if the system is linear-quadratic. But often this is not the case; solutions that assume linear dynamics and quadratic costs may fall apart when faced with nonlinear, realistic interactions, or when necessary convergence criteria are not met in Equation (11) [bacsar1998dynamic]. We therefore propose to use the solution from Section IV-A as a first pass approximation that we will refine with stochastic gradient descent.
As a reminder, our ultimate goal is to reach a robust robot policy. This policy is defined by Equation (5), which can be treated as a nested optimization problem. In the outer loop the robot proposes a sequence of actions $u^{0:T}$ to minimize its cost, and in the inner loop the antagonistic human responds with actions $w^{0:T}$ that maximize cost. We therefore apply a nested gradient descent algorithm to iteratively improve the initial action trajectory $\xi(x^{0})=(u^{0:T},w^{0:T})$ and optimize Equation (5). More specifically, we perform nested Metropolis-Hastings sampling (i.e., stochastic gradient descent). Other Monte Carlo methods can be leveraged within our general framework; we here selected the Metropolis-Hastings sampler due to its ergodic properties and avoidance of local optima [andrieu2006ergodicity]. Prior works suggest that LQ-approximate methods can be prone to these local optima [bacsar1998dynamic].
Given the seed $\xi(x^{0})=(u^{0:T},w^{0:T})$ provided by Section IV-A, this sampler has two parts [hoegerman2023reward]: an inner loop that updates the antagonistic human actions $\xi_{w}=w^{0:T}$, and an outer loop that improves the robot’s actions $\xi_{u}=u^{0:T}$. An outline of our resulting approach is included here.
@@
Summary. Taken together, the LQ game in Section IV-A provides a first guess at the robot’s action trajectory $\xi$. This action trajectory is then refined by nested Metropolis-Hastings samplers that update $\xi$ to solve for the true Nash Equilibrium policy in Equation (5). The resulting trajectory $\xi^{\prime}$ contains the sequence of actions a robust robot should execute over the next $T$ timesteps; this robust action sequence is recomputed online (i.e., at each timestep $t$).
We note that the metropolis-hastings sampler shown in Equations (15) and (16) are not guaranteed to output optimal trajectories. In fact, they may produce trajectories that exhibit higher cost than those produced by the DARE (Equation (11)) with probability (following from [chib1995understanding, tierney1994markov]):
$\begin{gathered}P(\|\xi^{\prime}-\xi^{\star}\|\leq r)\gtrsim 1-\mathcal{O}(e^{-\beta\Delta(r)})-\mathcal{O}(e^{-N})\\ \Delta(r)=\inf\limits_{\|\xi^{\prime}-\xi^{\star}\|\geq r}\left[J\left(\xi_{u}^{\prime},\xi_{w}^{\star}\right)-J\left(\xi_{u}^{\star},\xi_{w}^{\prime}\right)\right]\end{gathered}$ (17)
for a fixed but arbitrary $r$ after $N$ iterations. If $J$ is convex in $\xi_{u}$ and concave in $\xi_{w}$ (e.g., for quadratic cost functions like $J(x,u,w)=\|x\|_{Q}^{2}+\|u\|_{R_{u}}^{2}-\|w\|_{R_{w}}^{2}$), then:
$\begin{gathered}P\left(\|\xi^{\prime}-\xi^{\star}\|\leq r\right)\gtrsim 1-2e^{-\frac{\beta}{2}r^{2}}-\mathcal{O}(e^{-N})\end{gathered}$ (18)
So, for a higher $\beta$, the metropolis-hastings sampler is less likely to output trajectories that are strictly worse than the DARE. Note that the choice of seed also matters: if the DARE outputs optimal trajectories, it is probable that the metropolis hastings sampler’s refinement will be suboptimal. Now, we consider two cases for the system $\{f,J\}$. First, when the system has linear dynamics and quadratic cost, and second, when the system is not linear-quadratic.
@@
LQ Systems. For a linear-quadratic system the Bellman Equation is equivalent to the DARE presented in Equation (11) [zhang2021derivative]. Hence, the trajectory produced by the DARE is the Nash Equilibrium of the zero-sum game. This means that — for our MCLQ method — the initial trajectory is the global optimum of Equation (5), and Monte Carlo sampling is unable to improve upon this initial guess.
Non-LQ Systems. If the system is not linear-quadratic, then MCLQ improves upon the LQ approximation with a probability proportional to the error in the LQ approximation. As shown in Equation (17), if the error in the LQ approximation is very large, then the sampler will improve upon the initial trajectory with probability proportional to $1-\exp(-\beta\Delta)$.
IV-C Balancing Performance and Safety